BELLMAN FUNCTIONS AND TWO WEIGHT INEQUALITIES FOR HAAR MULTIPLIERS

FEDOR NAZAROV SERGEI TREIL ALEXANDER VOLBERG

Abstract. We give necessary and sufficient conditions for two weight norm inequalities for Haar multipliers operators and for square functions. We also give sufficient conditions for two weight norm inequalities for the Hilbert transform.